An engine absorbs 1.68 kJ from a hot reservoir at 277°C and expels 1.22 kJ to a cold reservoir at 27°C in each cycle.

(a) What is the engine's efficiency?
%

(b) How much work is done by the engine in each cycle?
J

(c) What is the power output of the engine if each cycle lasts 0.297 s?
kW
Answer #1

a) Engines efficiency, n= Work done heat added w o = On-ą Bu b) =1168- 1.22 )100 1.68 = 27.381 % Amount of work done by engi